# The underlying structures of self-attention: symmetry, directionality, and emergent dynamics in Transformer training

### v1

@@ 2.1. Interpreting self-attention with bilinear forms
Intuitively, this indicates that $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ define an alternative metric $\mathbf{W}_{qk}$ in the embedding space, which quantifies the score of $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ without requiring explicit query and key vectors. We formalize this intuition in the following,
@@
(Self-attention scores preserve the structure of geometrical projections). Let $\mathbf{X}\in\mathbb{R}^{N,d}$ be the embedding of a sequence of $N$ tokens such that $\mathbf{X}=[\mathbf{x}_{0}^{\top},\mathbf{x}_{1}^{\top},\dots,\mathbf{x}_{N}^{\top}]$ and $\mathbf{x}_{i}\in\mathbb{R}^{d}\,\,\forall\,i$. Let $\mathbf{A}$ be the matrix of attention scores, and let $\alpha_{ij}=[\mathbf{A}]_{ij}$ be the $(i,j)$ element of the attention matrix. We define $\mathbf{W}_{qk}=\mathbf{W}_{q}\mathbf{W}_{k}^{\top}$. The self-attention transformation maps the $i$-th token $\mathbf{x}_{i}$ into $\hat{\mathbf{x}}_{i}$ as follows,
$\hat{\mathbf{x}}^{\top}_{i}=\mathbf{x}^{\top}_{i}+\sum_{j=1}^{N}\alpha_{ij}\,\mathbf{x}_{j}\mathbf{W}_{v}\,,$ (3)
where the coefficients $\{\alpha_{ij}\}$ represent a convex combination of $\mathbf{x}_{j}$ within the subspace spanned by $\mathbf{X}=\{\mathbf{x}_{0}^{\top},\mathbf{x}_{1}^{\top},\dots,\mathbf{x}_{N}^{\top}\}$. These coefficients preserve the ordering of the geometrical projection of $\mathbf{x}_{i}$ onto the subspace $\mathbf{X}$ given by the bilinear form $\mathbf{W}_{qk}$,
$\hat{\alpha}_{ij}(\mathbf{W}_{qk})=\langle\mathbf{x}_{i},\mathbf{W}_{qk}\mathbf{x}_{j}\rangle=\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle_{\mathbf{W}_{qk}}\,,$ (4)
such that,
A proof of this proposition and its generalization for multi-head attention is in Appendix S1.2. The order of the attention scores is thereby preserved from the raw attention scores to the final attention scores (after soft-max).
It follows that $\mathbf{W}_{qk}$ defines the geometric structure that encodes the relevance of the token embedding $\mathbf{x}_{j}$ in predicting $\mathbf{x}_{i}$ through projections within the subspace (see Equation (3)). Therefore, token prediction depends on learning an optimal bilinear form $\mathbf{W}_{qk}$ to project tokens in the embedding space effectively. We note that $\mathbf{W}_{v}$ is a linear transformation that is applied independently to each projection in the sum and thus does not influence our derivation. In the following sections, we demonstrate how this equivalent formulation of self-attention provides a useful framework for analyzing the training of Transformer models. Specifically, we show that the choice of the objective function such as autoregressive prediction [Radford et al., 2019] or bidirectional training [Devlin et al., 2019] produces distinct structural patterns in $\mathbf{W}_{qk}$.
2.2. Deriving the gradients of self-attention with bilinear forms
Figure 1: Illustration of the computation of the self-attention score between token $\mathbf{x}_{i}$ and token $\mathbf{x}_{j}$ (forward pass, see Equation (3)), and its corresponding contribution to the weight update of $\mathbf{W}_{qk}$ (backward pass, see Equation (8)). The symbols $\oplus$”, $\otimes$”, and $\odot$” refer to the addition, multiplication, and dot product operations, respectively.
To show the connection between the objective function and the structures of self-attention matrices, we derive a convenient formulation for the weight update of $\mathbf{W}_{qk}$. We first formulate a sequence modeling problem with self-supervised training as follows. Let $U=\{t_{1},\dots,t_{N}\}$ be a sequence of $N$ tokens from a vocabulary of dimension $V$. Let $\mathcal{L}(U)$ be the negative log-likelihood of each token $t_{i}$, expressed as
$\mathcal{L}(U;\mathcal{W})=\sum_{i}\mathcal{L}(t_{i})=-\sum_{i}\log p(t_{i}\,|\,\{t_{j}:j\in\mathcal{C}_{i}\};\mathcal{W})\,,$ (6)
where $\mathcal{W}$ is the set of trainable parameters, and $\mathcal{C}_{i}\subset[0,1,\dots,N]$ is the set of indices defining the set of tokens $\{t_{j}\}$ of the conditional probability distribution. This allows us to isolate the contribution of each token $t_{j}$ to the prediction of the target token $t_{i}$. Here, we demonstrate that the updates to the matrix $\mathbf{W}_{qk}$ follow a structured pattern: the contribution of the token $t_{j}$ in the prediction of the embedding $t_{i}$ results in adding a rank-1 matrix $\mathbf{K}_{ij}$ to the matrix $\mathbf{W}_{qk}$. As a result, the total weight update to $\mathbf{W}_{qk}$ is expressed as a linear combination of these rank-1 matrices. We formalize this observation in the following proposition.
Proposition 2.2.
@@ 2.2. Deriving the gradients of self-attention with bilinear forms
(The implicit weight update as sum of rank-1 matrices). Let $U=\{t_{1},\dots,t_{N}\}$ be a sequence of $N$ tokens, and let $\mathcal{L}(U)$ be the negative log-likelihood
$\mathcal{L}(U;\mathcal{W})=\sum_{i}\mathcal{L}(t_{i})=-\sum_{i}\log p(t_{i}\,|\,\{t_{j}:j\in\mathcal{C}_{i}\};\mathcal{W})\,,$ (7)
with the conditional probability modeled with a Transformer model with learnable parameters $\mathcal{W}$. Let the self-attention function be defined with a bilinear form as in Equation (1). Following the gradient of $\mathcal{L}(U)$, the implicit weight update of $\mathbf{W}_{qk}^{l}$ at the $l$-th layer is proportional to the sum of the contributions of each $t_{j}\in\mathcal{C}_{i}$ in predicting $t_{i}$, as follows,
$\Delta\mathbf{W}_{qk}^{l}\propto\sum_{i}\sum_{j\in C_{i}}\Delta\mathbf{W}_{qk}^{l}\Big{|}_{t_{i}\leftarrow t_{j}}=\sum_{i}\sum_{j\in C_{i}}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}$ (8)
or equivalently to the sum of the contribution of each $t_{i}\in P_{j}$ when predicted by $t_{j}$,
$\Delta\mathbf{W}_{qk}^{l}\propto\sum_{i\in P_{j}}\sum_{j}\Delta\mathbf{W}_{qk}^{l}\Big{|}_{t_{i}\leftarrow t_{j}}=\sum_{i\in P_{j}}\sum_{j}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}\,,$ (9)
@@
where $\beta^{l}_{ij}$ is a scalar given by the contribution of the token embedding $\mathbf{x}_{j}$ to the error in predicting the token embedding $\mathbf{x}_{i}$ at the $l$-th self-attention layer, and where $\mathbf{K}^{l-1}_{ij}\in\mathbb{M}_{d}$ is a rank-1 matrix given by the outer product of the token embeddings at the previous layers $l-1$,
@@ 2.3. The relation between objective functions and structures in self-attention matrices
Finally, we show how the formulation of $\Delta\mathbf{W}_{qk}$ enables the analysis of the contribution of any given token $t^{*}$ to the weight updates, how this affects the properties of $\mathbf{W}_{qk}$, and how to relate these properties to the specific objective function. Indeed, Proposition 2.2 indicates that a token $t^{*}$ impacts the updates of $\mathbf{W}_{qk}$ differently when serving as context for predicting other tokens or being itself predicted. When a token $t^{*}$ serves as context ($t_{j}=t^{*}$), the embeddings of all predicted tokens contribute to the column space of $\mathbf{W}_{qk}$, where the update of the $k$-th column $\mathbf{w}_{\cdot,k}$ is given by
$\Delta\mathbf{w}_{\cdot,k}\bigg{|}_{t_{j}=t^{*}}=[\mathbf{x}_{t^{*}}]_{k}\left(\sum_{i\in P_{t}}\beta_{ij}\mathbf{x}_{i}\right)\,.$ (11)
Only the embedding of $t^{*}$ is instead added to the row space, where the update of the $m$-th row $\Delta\mathbf{w}_{m,\cdot}$ is given by
$\Delta\mathbf{w}_{m,\cdot}\bigg{|}_{t_{i}=t^{*}}=\left(\sum_{i\in P_{t}}\beta_{ij}[\mathbf{x}_{i}]_{k}\right)\mathbf{x}_{t^{*}}\,.$ (12)
Intuitively, using $t^{*}$ as context increases the dimensionality of the column space proportionally to the embeddings of the predicted tokens, while reducing the row space along the direction of the embedding of $t^{*}$. Conversely, when $t^{*}$ is being predicted ($t_{i}=t^{*}$), all token embeddings from the context are added to the row space of $\mathbf{W}_{qk}$, while only the embedding of $t^{*}$ is added to the column space. Importantly, any token $t^{*}$ is used differently as context or as prediction depending on the training objective. Autoregressive training implicitly introduces directionality by predicting each token based solely on its preceding tokens. In contrast, bidirectional training uses tokens as context and as predictions symmetrically. This fundamental difference between the two objective functions affects the weight updates of $\mathbf{W}_{qk}$, and in turn, the structures encoded in its columns and rows. We describe this in the following two informal Theorems.
@@
(Informal) (Autoregressive training leads to directional weight upates). Let $\Delta\mathbf{W}_{qk}$ be the weight update from Proposition 2.2, derived from an autoregressive objective function. There are, on average, more columns with high norms than rows with high norms, that is, there exist a $\gamma\in\mathbb{R}$ such that,
$\text{Pr}[||\mathbf{w}_{\cdot,k}||>w]>\text{Pr}[||\mathbf{w}_{m,\cdot}||>w]\,\,\ \forall\,w>\gamma\,,$ (13)
for any $k$-th column and $m$-th row, thereby inducing column dominance.
@@
(Informal) (Bidirectional training leads to symmetric weight upates). Let $U$ be an ordered sequence of $N$ tokens $U=[t_{1},\dots,t_{N}]$. Let $\Delta\mathbf{W}_{qk}$ be the weight update from Proposition 2.2, derived from a bidirectional objective function. Every pair $(i,j)$ with $i\neq j$ contributes to the weight update with a term that is approximately symmetric,
$\Delta\mathbf{W}_{qk}\big{|}_{t_{i}\leftrightarrow t_{j}}\approx\Delta\mathbf{W}_{qk}\big{|}^{\top}_{t_{i}\leftrightarrow t_{j}}\,,$ (14)
thereby inducing symmetric structures.
We provide a formal definition and proof of these two Theorems, and the related Propositions and Lemmas in Appendix S1.4.

### v2

@@ 2.1. Interpreting self-attention with bilinear forms
Intuitively, this equivalence shows that the matrices $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ together define an alternative metric in the embedding space, which quantifies the score of $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ without explicitly constructing the query and key vectors. Let us consider a specific input token $\mathbf{x}_{i}\in\mathbb{R}^{d}$. The self-attention layer maps it to an updated token $\hat{\mathbf{x}}_{i}\in\mathbb{R}^{d}$ as follows,
$\hat{\mathbf{x}}^{\top}_{i}=\mathbf{x}^{\top}_{i}+\sum_{j=1}^{N}\hat{\alpha}_{ij}\,\mathbf{x}_{j}\mathbf{W}_{v}\,,$ (3)
where the coefficients $\{\hat{\alpha}_{ij}\}$ in the second term represent the projection of $\mathbf{x}_{i}$ onto $\text{span}\{\mathbf{X}\}$ (the subspace spanned by $\mathbf{X}=\{\mathbf{x}_{0}^{\top},\mathbf{x}_{1}^{\top},\dots,\mathbf{x}_{N}^{\top}\})$ in the transformed embedding space defined by $\mathbf{W}_{qk}$,
$\sum_{j}\hat{\alpha}_{ij}\,\mathbf{x}_{j}=\sum_{j}\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle_{\mathbf{W}_{qk}}\,\mathbf{x}_{j}\,.$ (4)
Because the softmax function preserves the order of its input values, the ranking of the raw attention scores $\hat{\alpha}_{ij}$ is maintained in the final normalized scores $\alpha_{ij}$,
As a result, the attention weights ${\alpha_{ij}}$ define a convex combination of the token vectors ${\mathbf{x}_{j}}$, meaning the output lies within the convex hull of the input sequence $\text{Conv}(\mathbf{X})\subset\text{span}\{\mathbf{X}\}$ (see also Appendix S1.2). We note that $\mathbf{W}_{v}$ is a linear transformation that is applied independently to each projection in the sum and thus does not influence our derivation.
In the following sections, we demonstrate how this equivalent formulation of self-attention provides a useful framework for analyzing the training of Transformer models. Specifically, we show that the choice of the objective function such as autoregressive prediction [Radford et al., 2019] or bidirectional training [Devlin et al., 2019] produces distinct structural patterns in $\mathbf{W}_{qk}$.
2.2. Deriving the gradients of self-attention with bilinear forms
Figure 1: Illustration of the computation of the self-attention score between token $\mathbf{x}_{i}$ and token $\mathbf{x}_{j}$ (forward pass, see Equation (3)), and its corresponding contribution to the weight update of $\mathbf{W}_{qk}$ (backward pass, see Equation (8) and Equation (9)). The symbols $\oplus$”, $\otimes$”, and $\odot$” refer to the addition, multiplication, and dot product operations, respectively.
To show the connection between the objective function and the structures of self-attention matrices, we derive a convenient formulation for the weight update of $\mathbf{W}_{qk}$. We first formulate a sequence modeling problem with self-supervised training as follows. Let $U=\{t_{1},\dots,t_{N}\}$ be a sequence of $N$ tokens. For each position $i$, let $\mathcal{D}_{i}\subseteq\{1,2,\dots,N\}$ denote the conditioning set, that is, the set of indices corresponding to the tokens used to predict the token $t_{i}$. This formulation makes it possible to isolate and analyze the contribution of each token $t_{j}\in\mathcal{D}_{i}$ to the prediction of $t_{i}$. Let $\mathcal{L}(U)$ be the negative log-likelihood of each token $t_{i}$, expressed as,
$\mathcal{L}(U;\mathcal{W})=\sum_{i=1}^{N}\mathcal{L}(t_{i})=-\sum_{i=1}^{N}\log p_{\mathcal{W}}(t_{i}\,|\,t_{\mathcal{D}_{i}})\,,$ (6)
where $\mathcal{W}$ is the set of trainable parameters, and $p_{\mathcal{W}}(t_{i}\,|\,t_{\mathcal{D}_{i}})$ are the conditional densities parameterized by the model (see also Appendix S1.3).
Here, we demonstrate that the updates to the matrix $\mathbf{W}_{qk}$ follow a structured pattern: the contribution of the token $t_{j}$ in the prediction of the embedding $t_{i}$ results in adding a rank-1 matrix $\mathbf{K}_{ij}$ to the matrix $\mathbf{W}_{qk}$. As a result, the total weight update to $\mathbf{W}_{qk}$ is expressed as a linear combination of these rank-1 matrices. We formalize this observation in the following proposition.
@@ 2.2. Deriving the gradients of self-attention with bilinear forms
(The implicit weight update as sum of rank-1 matrices). Let $U=\{t_{1},\dots,t_{N}\}$ be a sequence of $N$ tokens, and let $\mathcal{L}(U;\mathcal{W})$ denote the negative log-likelihood of the sequence under a Transformer model with learnable parameters $\mathcal{W}$,
$\mathcal{L}(U;\mathcal{W})=\sum_{i}\mathcal{L}(t_{i})=-\sum_{i}\log p_{\mathcal{W}}(t_{i}\mid t_{\mathcal{D}_{i}})\,,$ (7)
where $\mathcal{D}_{i}\subseteq\{1,\dots,N\}$ is the set of indices used to condition the prediction of token $t_{i}$. Let the self-attention function be defined as in Equation (1). Following the gradient of $\mathcal{L}(U;\mathcal{W})$, the implicit gradient-based update to the bilinear weight matrix $\mathbf{W}_{qk}^{l}$ at the $l$-th self-attention layer is proportionally equivalent to:
  1. 1.
    The sum of contributions from each token $t_{j}:j\in\mathcal{C}_{i}$ used to predict $t_{i}$,
    $\Delta\mathbf{W}_{qk}^{l}\propto\sum_{i}\sum_{j\in\mathcal{C}_{i}}\Delta\mathbf{W}_{qk}^{l}\Big{|}_{t_{i}\leftarrow t_{j}}=\sum_{i}\sum_{j\in\mathcal{C}_{i}}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}\,,$ (8)
    where $\mathcal{C}_{i}$ denotes the set of context indices for predicting $t_{i}$.
  2. 2.
    The sum of contributions from each prediction target $t_{i}:i\in\mathcal{P}_{j}$ when predicted with $t_{j}$,
    $\Delta\mathbf{W}_{qk}^{l}\propto\sum_{i\in\mathcal{P}_{j}}\sum_{j}\Delta\mathbf{W}_{qk}^{l}\Big{|}_{t_{i}\leftarrow t_{j}}=\sum_{i\in\mathcal{P}_{j}}\sum_{j}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}\,,$ (9)
    where $\mathcal{P}_{j}$ denotes the set of tokens for which $t_{j}$ is included in the context.
@@
Here, $\beta^{l}_{ij}$ is a scalar that quantifies the contribution of the token embedding $\mathbf{x}_{j}$ to the prediction error for $\mathbf{x}_{i}$ at the $l$-th layer, and the matrix $\mathbf{K}^{l-1}_{ij}\in\mathbb{M}_{d}$ is a rank-1 matrix defined by the outer product of the token embeddings at the previous layer,
@@
2.3. How context and prediction impact the gradient differently
Next, we show how the formulation of $\Delta\mathbf{W}_{qk}$ enables the analysis of the contribution of any given token $t^{*}$ to the weight updates and how this affects the properties of $\mathbf{W}_{qk}$. Indeed, Proposition 2.1 indicates that a token $t^{*}$ impacts the updates of $\mathbf{W}_{qk}$ differently when serving as context for predicting other tokens or being itself predicted.
When a token $t^{*}$ serves as context ($t_{j}=t^{*}$), the embeddings of all predicted tokens contribute to the column space of $\mathbf{W}_{qk}$, where the update of every $k$-th column $\mathbf{w}_{\cdot,k}$ is given by
$\Delta\mathbf{w}_{\cdot,k}\Big{|}_{t_{j}=t^{*}}=[\mathbf{x}_{t^{*}}]_{k}\left(\sum_{i\in\mathcal{P}_{t}}\beta_{i*}\mathbf{x}_{i}\right)\,.$ (11)
Only the embedding of $t^{*}$ is instead added to the row space, where the update of every $m$-th row $\Delta\mathbf{w}_{m,\cdot}$ is given by
$\Delta\mathbf{w}_{m,\cdot}\Big{|}_{t_{i}=t^{*}}=\left(\sum_{i\in\mathcal{P}_{t}}\beta_{i*}[\mathbf{x}_{i}]_{k}\right)\mathbf{x}_{t^{*}}\,.$ (12)
Intuitively, using $t^{*}$ as context increases the dimensionality of the column space proportionally to the embeddings of the predicted tokens, while reducing the row space along the direction of the embedding of $t^{*}$. Conversely, when $t^{*}$ is being predicted ($t_{i}=t^{*}$), all token embeddings from the context are added to the row space of $\mathbf{W}_{qk}$, while only the embedding of $t^{*}$ is added to the column space. Consequently, the role a token plays during training affects its contribution to the weight update differently. We formalize this in the following proposition.
Proposition 2.2.
(Different implicit updates for context and prediction). Let $U=\{t_{1},\dots,t_{N}\}$ be a sequence of tokens and let $\mathbf{x}_{i}\in\mathbb{R}^{d}$ be the embedding of the $i$-th token. Let $\Delta\mathbf{W}_{qk}$ be the weight update from Proposition 2.1. Let $t^{*}$ be a given token in $U$. When using $t^{*}$ as context, the $k$-th column of $\Delta\mathbf{W}_{qk}$ is given by
$\Delta\mathbf{w}_{\cdot,k}\bigg{|}_{t_{j}=t^{*}}=[\mathbf{x}_{t^{*}}]_{k}\left(\sum_{i\in\mathcal{P}_{t^{*}}}\beta_{i*}\mathbf{x}_{i}\right)\,,$ (13)
while the $m$-th row of $\Delta\mathbf{W}_{qk}$ is given by
$\Delta\mathbf{w}_{m,\cdot}\bigg{|}_{t_{j}=t^{*}}=\left(\sum_{i\in\mathcal{P}_{t^{*}}}\beta_{i*}[\mathbf{x}_{i}]_{m}\right)\mathbf{x}_{t^{*}}\,.$ (14)
When predicting $t^{*}$, the $k$-th column of $\Delta\mathbf{W}_{qk}$ is given by
$\Delta\mathbf{w}_{k,\cdot}\bigg{|}_{t_{i}=t^{*}}=\left(\sum_{j\in\mathcal{C}_{t^{*}}}\beta_{*j}[\mathbf{x}_{j}]_{k}\right)\mathbf{x}_{t^{*}}\,,$ (15)
while the $m$-th row of $\Delta\mathbf{W}_{qk}$ is given by
$\Delta\mathbf{w}_{m,\cdot}\bigg{|}_{t_{i}=t^{*}}=[\mathbf{x}_{t^{*}}]_{m}\left(\sum_{j\in\mathcal{C}_{t^{*}}}\beta_{*j}\mathbf{x}_{j}^{\top}\right)\,.$ (16)
We provide proof for this proposition with related remarks in Appendix S1.5.1. Note that all mathematical results derived so far rely solely on the structure of self-attention and make no assumptions about the input data. In the next section, we build on these results to link structural patterns in $\mathbf{W}_{qk}$ with the specific form of the objective function.
@@ 2.4. The relation between objective functions and structures in self-attention matrices
In this final section, we show how to relate these properties to the specific objective function. Crucially, the number of times a token $t^{*}$ appears as context or as a prediction depends on the training objective. Autoregressive training implicitly introduces directionality by predicting each token solely on its preceding tokens. In contrast, bidirectional training uses tokens as context and as predictions symmetrically. This fundamental difference affects the weight updates of $\mathbf{W}_{qk}$, and consequently, the structures encoded in its rows and columns.
We formalize this relationship in the next two theorems, under the following assumptions: (a) tokens exhibit statistical correlations, leading to partial alignment in their embeddings; (b) the entries of $\mathbf{W}_{qk}$ are i.i.d. at initialization with finite mean and variance; (c) some tokens tend to occur earlier in the sequence and are more predictive of future content; (d) bidirectional training induces approximately symmetric error signals. In this setting, we prove how the objective function influences the internal structure encoded by self-attention.
@@
(Autoregressive training induces directionality) Let $\mathcal{V}=\{t_{0},\dots,t_{V}\}$ be a vocabulary of tokens, and let $\mathcal{U}\subset\mathcal{V}^{N}$ denote the sample space of all sequences of length $N$. Let $U=\{t_{1},\dots,t_{N}\}\in\mathcal{U}$ be a random variable with $U\sim P(U)$. Define $\Pr[t_{j}=t^{*}]$ as the marginal probability that the token at position $j$ equals $t^{*}\in\mathcal{V}$. Let $\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ be the token embeddings corresponding to the elements of $U$, where each embedding $\mathbf{x}_{i}\sim\mathcal{D}$ is drawn i.i.d. from a distribution $\mathcal{D}$ with zero mean and covariance matrix $\text{Cov}(\mathbf{x}_{i})=\Sigma$.. Let $\mathbf{W}_{qk}$ be query-key matrix of a self-attention mechanism, and let $\Delta\mathbf{W}_{qk}$ denote its gradient update as defined in Proposition 2.1, computed under an autoregressive objective as in Definition S1.3.
It follows that the contribution of the token $t^{*}$ after training satisfies,
$\frac{\mathbb{E}_{\mathcal{D}}\left[\|\Delta\mathbf{w}_{\cdot,k}\|^{2}\right]}{\mathbb{E}_{\mathcal{D}}\left[\|\Delta\mathbf{w}_{m,\cdot}\|^{2}\right]}>1\quad\forall k,\,\,\forall m\,\,\text{s.t.}\,\,\Sigma_{m,m}<\frac{\mathrm{Tr}(\Sigma)}{d}\,.$ (17)
Moreover, there exists a scalar $\gamma\in\mathbb{R}_{>0}$, proportional to the product of the row and column norm variances, $\gamma\propto\mathrm{Var}\left(\|\mathbf{w}_{m,\cdot}\|\right)\cdot\mathrm{Var}\left(\|\mathbf{w}_{\cdot,k}\|\right)$, such that for all $w>\gamma$,
$\Pr\left[\|\mathbf{w}_{\cdot,k}\|^{2}>w\right]>\Pr\left[\|\mathbf{w}_{m,\cdot}\|^{2}>w\right]\,,$ (18)
that is, columns of $\mathbf{W}_{qk}$ are more likely than rows to have high norm under autoregressive training, indicating a directional bias in gradient updates.
@@
(Bidirectional training induces symmetry). Let $U=\{t_{1},\dots,t_{N}\}$ be a sequence of tokens. Let $\mathbf{W}_{qk}$ be the query-key matrix of a self-attention mechanism, and let $\Delta\mathbf{W}_{qk}$ denote its gradient update as defined in Proposition 2.1, computed under a bidirectional objective as in Definition S1.3. It follows that,
$\Delta\mathbf{W}_{qk}=\sum_{i=1}^{N}\sum_{j=1}^{N}\beta_{ij}\mathbf{K}_{ij}\,,$ (19)
$\Delta\mathbf{W}_{qk}\big{|}_{\mathbf{t}_{i}\leftrightarrow\mathbf{t}_{j}}=\beta_{ij}\mathbf{K}_{ij}+\beta_{ji}{\mathbf{K}_{ij}}^{T}\,,$ (20)
that is approximately symmetric,
$\Delta\mathbf{W}_{qk}\big{|}_{\mathbf{t}_{i}\leftrightarrow\mathbf{t}_{j}}\approx\Delta\mathbf{W}_{qk}\big{|}^{\top}_{\mathbf{t}_{i}\leftrightarrow\mathbf{t}_{j}}\,.$ (21)
We provide a proof of these two Theorems, and the related Propositions and Lemmas in Appendix S1.5.